OpenStudy (anonymous):

is this series geometric -1,1,4,8

OpenStudy (amistre64):

hmm, a series is a sum of the terms of a sequence, not the sequence itself .... what is your definition of a geometric sequence?

OpenStudy (anonymous):

A common ratio to describe how to get from on to the next

OpenStudy (amistre64):

then lets see if we can find a common ratio -1 * r = 1 1 * r = 4 4 * r = 8 hmm, does this look doable to you?

OpenStudy (anonymous):

Not too sure I understand

OpenStudy (amistre64):

well in order to get from the first term of 1 to the second term of -1, the for some value of r, we just multiply it out same idea for 2nd to 3rd, and 3rd to 4th

OpenStudy (amistre64):

if we solve each instance for r, and find that it is the same each time, then there is a common r (for ratio) that gets us from one term to the next

OpenStudy (anonymous):

It looks like a consecutive integer problem in a way

OpenStudy (amistre64):

almost :) but lets solve each instance for r to see if it remains the same -1 * r = 1 r = -1 1 * r = 4 r = 4 4 * r = 8 r = 2 since r is changing, the ratio between terms is not constant .... so this cant be geometric

OpenStudy (anonymous):

Ok thank you so much

OpenStudy (amistre64):

yep